Schedule of Cash Payments

Excel Learning Systems Inc. was organized on September 30, 2016. Projected selling and administrative expenses for each of the first three months of operations are as follows:

October $109,000
November 103,600
December 94,300

Depreciation, insurance, and property taxes represent $24,000 of the estimated monthly expenses. The annual insurance premium was paid on September 30, and property taxes for the year will be paid in June. The company expects that 58% of the remainder of the expenses will be paid in the month in which they are incurred, with the balance to be paid in the following month.

Prepare a schedule indicating cash payments for selling and administrative expenses for October, November, and December.

Excel Learning Systems Inc.
Schedule of Cash Payments for Selling and Administrative Expenses
For the Three Months Ending December 31, 2016
October November December
October expenses:
Paid in October $
Paid in November $
November expenses:
Paid in November
Paid in December $
December expenses:
Paid in December
Total cash payments $ $ $
Answer #1
Excel Learning Systems Inc.
Schedule of Cash Payments for Selling and Administrative Expenses
For the Three Months Ending December 31, 2016
October November December
October expenses:
Paid in October 49300
Paid in November 35700
November expenses:
Paid in November 46168
Paid in December 33432
December expenses:
Paid in December 40774
Total cash payments 49300 81868 74206
Workings:
October November December
October expenses:
Paid in October =(109000-24000)*58%
Paid in November =(109000-24000)*42%
November expenses:
Paid in November =(103600-24000)*58%
Paid in December =(103600-24000)*42%
December expenses:
Paid in December =(94300-24000)*58%